b. At prices of P1 = $5 and P2 = $6 quantities demanded are Q1 = 30 - 2*5 = 20, Q2 = 30 - 2*6 = 18, so the arc elasticity is:

Ed = (18 - 20)/(6 - 5)*(6 + 5)/(18 + 20) = -2/1*11/38 = -0.579. please give me formula because arc elasticity formula is different
1
Expert's answer
2020-03-10T10:55:33-0400

b. At prices of P1 = $5 and P2 = $6 quantities demanded are Q1 = 30 - 2*5 = 20, Q2 = 30 - 2*6 = 18, so the arc elasticity is:

"Ed = \frac{Q2 - Q1}{(Q2 + Q1)/2}×\frac{(P2 + P1)/2}{P2 - P1} = \frac{18 - 20}{(18 + 20)/2}×\frac{(6 + 5)/2}{6 - 5} = -2/19×5.5/1 = -0.579."